OpenStudy (anonymous):

Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring? −b, b2 − 4ac, 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (anonymous):

b^2 - 4ac, it is the discriminant, and if it is 0+ it can be factored. If it is less than 0, it cannot be factored.

OpenStudy (anonymous):

(b^2 - 4ac)^1/2 tells you for the first part unless you are using complex

OpenStudy (anonymous):

If the discriminant is a perfect square, the quadratic equation is factorable. I've never seen this written in a book, but one of my professors taught it to me.

OpenStudy (inkyvoyd):

Well, that's obvious

OpenStudy (inkyvoyd):

you have to take the square root of what's in the discriminant

OpenStudy (anonymous):

answer to the question is (2x+1)(x+3).

OpenStudy (inkyvoyd):

in fact, that's the only part of the root that could be irrational

OpenStudy (inkyvoyd):

And wiz, don't do that.

OpenStudy (inkyvoyd):

@wizguy

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i did this stuff in grade 5 inkyvoyd, so dont do what?

OpenStudy (inkyvoyd):

Don't give out the answer.

OpenStudy (inkyvoyd):

It's against the CoC to give it out directly. You didn't do that, but it's not exactly in the spirit of OS to give out the answer like that.

OpenStudy (anonymous):

by any chance is it 7^2-4*2*3=49-24=25=5^2 o nevermind I thought that wizguy just gave the phony answer

OpenStudy (inkyvoyd):

You just basically told someone how to find the derivative of a polynomial, then gave them the answer to the question

OpenStudy (inkyvoyd):

without walking them through hwo to find the "derivative of a polynomial"

OpenStudy (anonymous):

but thank you guys once again

OpenStudy (inkyvoyd):

That's techinically looked down upon, and it's close to answer sniping.